c. Calculate the confidence level that would be required to have the mean



assembly time estimated up to ±15 seconds assuming the sample size



is 120 minutes.

1
Expert's answer
2021-12-10T08:52:37-0500

Given "n=120, E=15\ s=0.25\ min"


"E=z_c\times\dfrac{\sigma}{\sqrt{n}}"

"z_c=\dfrac{E(\sqrt{n})}{\sigma}"

"z_c=\dfrac{0.25(\sqrt{120})}{\sigma}"

Let "\sigma=2\ min"


"z_c=\dfrac{0.25(\sqrt{120})}{2}\approx1.3693"

"P(Z>1.3693)=0.08545"

"2(0.08545)=0.1709"

"1-0.1709=0.8291"

"82.91"% confidence interval